# Towards Multi-modal Transformers in Federated Learning

### v1

@@ 0.A.4 Distribution of the data
Figure 7: Multi-model FL in the vision-language domain with collaboration from different modalities.
@@ 0.B.1 Implementation Details
We use AdamW [33] as the optimizer with a learning rate of $0.0001$ with a decay of $0.99$ every epoch for local training. The batch size is set as $112$ in most cases. All the experiments are implemented under the PyTorch framework and run on 4×4$\times$4 × Nvidia A5000 GPUs. We will release our codes after the paper is accepted.
@@ 0.B.5 Imbalanced client scenario
0.B.6 Communication analysis
Table 8: Communication cost and performance of each method on Flickr
Method Comm. Cost (MB) $\mathbf{R@1_{\text{sum}}}$
FedAvg $208.81$ $81.08$
FedProx $208.81$ $78.55$
CreamFL $211.74$ $74.83$
FedIoT $208.81$ $85.51$
FedCola (CA-only) $208.81$ $90.09$
FedCola (Attn) $262.95$ $91.73$
FedCola $371.26$ $91.96$

### v2

@@ 0.A.4 Distribution of the data
Figure 6: Multi-modal FL in the vision-language domain with collaboration from different modalities.
@@ 0.B.1 Implementation Details
We use AdamW [36] as the optimizer with a learning rate of $0.0001$ with a decay of $0.99$ every epoch for local training. The batch size is set as $112$ in most cases. All the experiments are implemented under the PyTorch framework and run on 4×4$\times$4 × Nvidia A5000 GPUs.
@@ 0.B.3 Stochasticity discussion
0.B.4 Breakdown performance
0.B.5 Communication analysis
Table 6: Communication cost and performance of each method on Flickr
Method Comm. Cost (MB) $\mathbf{R@1_{\text{sum}}}$
FedAvg $208.81$ $81.08$
FedProx $208.81$ $78.55$
CreamFL $211.74$ $74.83$
FedIoT $208.81$ $85.51$
FedCola (CA-only) $208.81$ $90.09$
FedCola (Attn) $262.95$ $91.73$
FedCola $371.26$ $91.96$
0.B.6 Imbalanced client scenario